Mission Operations Days: 2004/115 to 2004/121
April 24, 2004 00:00:00 GMT to April 30, 2004 00:00:00 GMT
Terra Spacecraft and MODIS Instrument Status:
|AM-1 "Terra" is in Normal Mode|
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
|Blackbody||A On (290K); B Off||Nominal Science Mode|
|Calibration Electronics||A On; B Off||Nominal Science Mode|
|Control Processor||A On; B Off||Nominal Science Mode|
|Nadir Door||Unlatched, Open||Nominal Science Mode|
|Space View Door||Unlatched, Open||Nominal Science Mode|
|Solar Diffuser Door||Unlatched, Screened||Permanent Configuration|
|FDDI Formatter||A On; B Off||Nominal Science Mode|
|FDDI Port||A On; B Off||Nominal Science Mode|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal Science Mode|
|Format Processor||A Off; B On||Nominal Science Mode|
|Power Supplies||#1 On, #2 Off||Nominal Science Mode|
|PV VIS FPA||A On; B Off||Nominal Science Mode|
|PV NIR FPA||A On; B Off||Nominal Science Mode|
|PV SMIR FPA||A On; B Off||Nominal Science Mode|
|PV LWIR FPA||A On; B Off||Nominal Science Mode|
|PC LWIR FPA||A On; B Off||Nominal Science Mode|
|RC Outgas Heaters||All 3 (CS, IS, OS) Off||Nominal Science Mode|
|RC LWIR FPA Heater||On (83K)||Nominal Science Mode|
|RC SMIR FPA Heater||Off||Nominal Science Mode|
|Scan Assembly||A On; B Off||Nominal Science Mode|
|SDSM||Off||Nominal Science Mode|
|SRCA||Off||Nominal Science Mode|
|PS1 Survival Heater||Enabled||Nominal Science Mode|
|PS2 Survival Heater||Enabled||Nominal Science Mode|
|Timing Generator||A On; B Off||Nominal Science Mode|
|Flight Software||Rev BD + 11 sets of patches||Nominal Science Mode|
|Inhibit Ids Set||none||Nominal Science Mode|
|TMON 66||Disabled||Nominal Science Mode|
|TMON 67||Disabled||Nominal Science Mode|
|TMON 68||Enabled||Nominal Science Mode|
|TMON 69||Disabled||Nominal Science Mode|
This Week's Completed MODIS Activities:
Saturday, April 24, 2004
2004/115 00:00:00 ATC MOD_OA26_BB_285K_A.1 Set blackbody temperature to 285K on the A Side
2004/115 03:00:00 ATC MOD_OA26_BB_290K_A.1 Set blackbody temperature to 290K on the A Side
2004/115 07:00:00 ATC MOD_OA26_BB_315K_A.1 Set blackbody temperature to 315K on the A Side
2004/115 07:00:30 ATC MOD_OA26_BB_DCYCLE_FULL.1 Set blackbody heater duty cycle to FULL
2004/115 10:00:00 ATC MOD_OA26_BB_270K_A.1 Set blackbody temperature to 270K on the A Side
2004/115 10:00:30 ATC MOD_OA26_BB_DCYCLE_THIRD.1 Set blackbody heater duty cycle to THIRD
Sunday, April 25, 2004
2004/116 07:00:00 ATC MOD_OA26_BB_290K_A.1 Set blackbody temperature to 290K on the A Side
Monday, April 26, 2004
2004/117 09:00:00 ATC MOD_SRCA_14_6_4_2_SETUP.3 Setup of non-nominal lamp configuration for SRCA calibration
2004/117 09:13:48-09:34:43 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #2 Data collection from 09:29:42-09:34:43.
2004/117 14:00:00 ATC MOD_SRCA_14_6_2_2_SETUP.3 Setup of non-nominal lamp configuration for SRCA calibration
2004/117 14:10:26-14:31:21 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #3 Data collection from 14:26:20-14:31:21.
Tuesday, April 27, 2004
2004/118 20:39:51-20:48:39 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
2004/118 21:40:00-21:40:31 ATC MOD_SRCA_NOMINAL_SETUP.4 Setup of nominal lamp configuration for SRCA calibration
2004/118 21:43:42-22:04:36 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #1 Data Collect (21:59:36-22:04:36) Centered on S/C Night
Wednesday, April 28, 2004
Thursday, April 29, 2004
Friday, April 30, 2004
This Week's Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Calibrations for May will include weekly SDSM, monthly SRCA Full and 10W Radiometric, and bimonthly SRCA Full Spatial calibrations.
There will be no Lunar Calibration Rolls in April, May, or June for the Terra MODIS. The necessary roll angles are -26.3051, -27.2921, and -22.1398 degrees, respectively; all of which are beyond the -20.0 degree limit.
The next DMU for Terra is expected to occur on June 1, 2004. This date has been chosen to avoid the Aura Launch and the ASTER AOS Server replacement. Details on the duration and time of the maneuver have not yet been determined.
MODIS NAD (Nadir Aperture Door) Anomaly:
Telemetry still indicates that the “NAD Open” (MOD_CR_DR_NAD_OPEN) trigger is set to the “NOT_OPEN” position; however, science data has shown that the NAD is open. The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.
General Instrument Comments:
MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.
MODIS Telemetry Trends:
Nominal telemetry trending is ongoing.
Non-MODIS Significant Events:
Terra Solid State Recorder Anomaly
The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA's that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous reports.
The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.
Limited Life Item Status:
SRCA 10W Lamp #1: 230.1 of 500 hours
SRCA 10W Lamp #2: 168.7 of 500 hours
SRCA 10W Lamp #3: 177.9 of 500 hours
SRCA 10W Lamp #4: 62.6 of 500 hours
SRCA 1W Lamp #1: 565.9 of 4000 hours
SRCA 1W Lamp #2: 276.9 of 4000 hours
Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements